# Regime Learning for Differentiable Particle Filters

### v2

@@ I Introduction
In this paper, we introduce the regime learning particle filter, which uses neural networks to parameterise the switching dynamic. Our contributions are two-fold:
@@
The structure of the paper is as follows: in Section II we formally set-up the problem the paper addresses; in Section III we review the required background knowledge. Section IV introduces the proposed regime learning particle filter algorithm, with Sections IV-A and IV-B introducing a novel parameterisation and Section IV-C devoted to a proposed training strategy. In Section V we demonstrate competitive performance, with respect to prior state-of-the-art approaches, on the test environments adopted in previous works. We conclude this paper in Section VI.
@@ III-A Regime switching particle filtering
The mean, $\hat{\mathbf{x}}_{t}$, relies on auto-normalised importance weights so is biased in general [12], but $\hat{p}\left({\mathbf{y}_{0:t}}\right)$ is unbiased [13]. The derivation of both equations can be found in [12].
@@ III-B Differentiable particle filtering
The regime switching differentiable particle filter (RSDBPF) was proposed in [9] to address the regime-switching filtering problem (LABEL:set-up) where the switching dynamic is assumed to be known but the other model components are not. It can be thought of a RSPF where the parameters are learned by gradient descent. They choose not to employ soft-resampling, and instead zero out the contribution to the gradient of the particles due their ancestors at every time-step, improving gradient variance at the cost of bias. This is found empirically to improve performance in some cases [20].
@@ IV-B Parameterising the switching dynamic
We propose a neural network parameterisation of the switching dynamic. It is well known that particle filters for which the late-time state depends strongly on the early-time state do not preform well, in fact, permitting strong dependence, one can construct filters that diverge in variance at any desired rate [12]. Late-time particles are genealogically non-diverse at early time and so form poor samples of the early-time state; a phenomenon often referred to as “path-degeneracy”. For this reason, we take inspiration from the architecture of the long-short term memory (LSTM) unit [21] and build forget gates into our model as:
@@ IV-C Training the RLPF
We adopt a novel training strategy. Instead of treating the problem as a supervised regression on $\hat{\mathbf{x}}_{0:T}$ with the MSE-loss $\mathcal{L}_{\text{MSE}}\left({\hat{\mathbf{x}}_{0:T},\mathbf{x}_{0:T}}\right)$ (Eq. (4)) as in previous work [9], we introduce an additional $\mathcal{L}_{\text{ELBO}}\left({\left\{{\mathbf{x}_{0:T},\mathbf{y}_{0:T}}\right\}}\right)$ (Eq. (5)) unsupervised loss term in the training objective.

### v3

@@ I Introduction
00footnotetext: The authors acknowledge funding from the UK Department for Science, Innovation and Technology through the 2024 National Measurement System programme. JJ. Brady is supported through the National Physical Laboratory and University of Surrey partnership via an Engineering and Physical Sciences Research Council studentship. The work of V. E. is supported by ARL/ARO under grant W911NF-22-1-0235 and the University of Edinburgh – Rice University Strategic Collaboration Award.
@@
In this paper, we introduce the regime learning particle filter (RLPF), which uses neural networks to parameterise the switching dynamic. Our contributions are two-fold:
@@
The structure of the paper is as follows: in Section II we formally set up the problem the paper addresses; in Section III we review the required background knowledge. Section IV introduces the proposed regime learning particle filter algorithm, with Sections IV-A and IV-B introducing a novel parameterisation and Section IV-C devoted to a proposed training strategy. In Section V we demonstrate competitive performance, with respect to prior state-of-the-art approaches, on the test environments adopted in previous works. We conclude this paper in Section VI.
@@ III-A Regime switching particle filtering
The mean, $\hat{\mathbf{x}}_{t}$, relies on auto-normalised importance weights and so is biased in general [12], but $\hat{p}\left({\mathbf{y}_{0:t}}\right)$ is unbiased [13]. The derivation of both equations can be found in [12].
@@ III-B Differentiable particle filtering
The regime switching differentiable particle filter (RSDBPF) was proposed in [9] to address the regime-switching filtering problem (LABEL:set-up) where the switching dynamic is assumed to be known but the other model components are not. It can be thought of a RSPF where the parameters are learned by gradient descent. The RSDBPF does not employ soft-resampling, and instead zeros out the contribution to the gradient of the particles due their ancestors at every time-step, improving gradient variance at the cost of bias. This is found empirically to improve performance in some cases [20].
@@ IV-B Parameterising the switching dynamic
We propose a neural network parameterisation of the switching dynamic. It is well known that particle filters for which the late-time state depends strongly on the early-time state do not perform well, in fact, permitting strong dependence, one can construct filters that diverge in variance at any desired rate [12]. Late-time particles are genealogically non-diverse at early time and so form poor samples of the early-time state; a phenomenon often referred to as “path-degeneracy”. For this reason, we take inspiration from the architecture of the long-short term memory (LSTM) unit [21] and build forget gates into our model as:
@@ IV-C Training the RLPF
We adopt a novel training strategy. Instead of treating the problem as a supervised regression on $\hat{\mathbf{x}}_{0:T}$ with the MSE-loss $\mathcal{L}_{\text{MSE}}\left({\hat{\mathbf{x}}_{0:T},\mathbf{x}_{0:T}}\right)$ (Eq. (4)) as in previous work [9], we introduce an additional $\mathcal{L}_{\text{ELBO}}\left({\left\{{\mathbf{x}_{0:T},\mathbf{y}_{0:T}}\right\}}\right)$ (Eq. (5)) unsupervised loss term in the training objective: